OpenStudy (anonymous):

Consider the following reaction, which is carried out at 300 K: CH3CH2OH + 3O2 → 2CO2 + 3H2O ∆H = –1,235 kJ/mol and ∆S = +0.22kJ/mol•K • What is the value of ∆G? Show your work. • Classify the reaction as spontaneous or nonspontaneous. • Suggest why the value for ∆S is positive in this reaction.

OpenStudy (frostbite):

1) \[\Large \Delta G = \Delta H - T \Delta S\] 2) Is the \(\Delta G\) negative or positive? 3) Any good suggestion what might be the answer? try look at how much is being created from how much is being used. [Entropy being a measure for order and disorder OR the distribution of stuff]

OpenStudy (frostbite):

We can go over each one if you like, but I want you to solve the problem.

OpenStudy (anonymous):

yes please

OpenStudy (frostbite):

Okay. Lets look at the first one. The most known equation for Gibbs free energy at constant pressure at temperature \(T\) is\[\Large \Delta G = \Delta H- T \Delta S\] Here are, \(\Delta G\) change in Gibbs free energy, \(\Delta H\) the change in enthalpy and \(\Delta S\) the change in entropy. As the information we have been give is all at the same temperature we can calculate the change in Gibbs free energy at 300 K. Try plug in the numbers given in the equation above and calculate the Gibbs free energy. (I check your answer when you post it)

OpenStudy (anonymous):

so so i plug in 1.235 and .22 to the formual ?

OpenStudy (frostbite):

Yes and the 300 at T

OpenStudy (anonymous):

I got 205.7

OpenStudy (frostbite):

Hmmm I think you have done something wrong. Did you calculate the following? \[\Large \Delta G=-1235 \frac{ kJ }{ mol }-300 ~ K \times 0.22 \frac{ kJ }{ mol \times K }\]

OpenStudy (frostbite):

As you can see the overall result is going to be negative

OpenStudy (anonymous):

Oh I did know it was negative. But this time I got -357.5

OpenStudy (frostbite):

I don't see how you get that number. Lets break it up by parts: We got from above that: \[\Large \Delta G=-1235 \frac{ kJ }{ mol }-300 ~ K \times 0.22 \frac{ kJ }{ mol \times K }\] And notice that \[\Large \Delta H =-1235 \frac{ kJ }{ mol }\] \[\Large T \Delta S=300 K \times 0.22 \frac{ kJ }{ K \times mol }= 66\frac{ kJ }{ mol }\] \[\Large \Delta G=-1235 \frac{ kJ }{ mol }-66 \frac{ kJ }{ mol }=1301 \frac{ kJ }{ mol }\] @aaronq please check... I'm getting tired and want to be sure I don't make a mistake.

OpenStudy (frostbite):

Correction! \[\Large \Delta G=-1301 \frac{ kJ }{ mol }\]

OpenStudy (anonymous):

Oh okay i was mutliplying wrong

OpenStudy (frostbite):

It happens :) So the answer is (Just want you to write it)

OpenStudy (anonymous):

-1301kJ/mol

OpenStudy (frostbite):

Exactly. :) that is the answer to 1)

OpenStudy (frostbite):

For number 2 we are going to use some knowledge about the Gibbs free energy. \(\Delta G\) is a among others a indicator for spontaneous change. The ability for the reaction to happen spontaneously is given weather \(\Delta G\) is positive or negative. Do you know, or can you guess, if \(\Delta G\) should be positive or negative for a spontaneous reaction to occur?

OpenStudy (anonymous):

I think it should be negative

OpenStudy (frostbite):

You are thinking correctly! For a spontaneous reaction \(\Delta G<0\), and for a non-spontaneous reaction \(\Delta G>0\). So is the reaction CH3CH2OH + 3O2 → 2CO2 + 3H2O spontaneous at 300 K?

OpenStudy (anonymous):

its spontaneous right?

OpenStudy (frostbite):

It sure is. Well done so far! :) Now to the last one, you need to know what entropy (\(\Delta S\)) is. Do you have an idea what entropy is about?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I get it

OpenStudy (frostbite):

Right sorry I deleted the answer, but I try make a general conclusion you should just remember (it will work until you study thermodynamics more closely): "For a change from a state we call X into a state we call Y (donated \(X \to Y\)), the state with the most accessible energy levels / most accessible states, will be favorer by an increase in entropy."

OpenStudy (frostbite):

So for your example it would be: CH3CH2OH + 3O2 → 2CO2 + 3H2O Note we got 4 particles on the reactant side, but 5 particles on the product side. Using the general conclusion we can write \(4 X \to 5Y\) Meaning we got a positive entropy.

OpenStudy (anonymous):

So we are trying to change it from a negative to a positive?

OpenStudy (frostbite):

Not really. Entropy is all about how many accessible states particles can take (vibrational, translational, rotational ect.). But for simplicity we ONLY focus on how many particles there are on each side of the reaction and assume that all particles contributes with exactly one state.

OpenStudy (frostbite):